OpenStudy (anonymous):

If salary of A is greater than B by 50% then salary of B is how much % less than A?

OpenStudy (anonymous):

Equations: A = 1.5B Solve for "B":: B = A/(1.5) = A/(3/2) = (2/3)A

OpenStudy (anonymous):

B gets \[\frac{ 1 }{ 3 } \]

OpenStudy (anonymous):

or 11% less than A

OpenStudy (dm91):

ayeshaafzal is right it would be 1/3 or 11% less than A

OpenStudy (rhystic):

how did 1/3 = 11%?

OpenStudy (rhystic):

The math is correct... the conversion of the answer (fraction) to (percentage) is wrong.....

OpenStudy (anonymous):

No it's not 11% but how did you got 1/3. The ans is (1/3)100=33.33%

OpenStudy (rhystic):

So you have A which is 50% greater than B. Converting that to a number from percentage you get 1.50 (because its 50 % greater meaning 150% ... bring that to a decimal you shift to the left two decimal points yielding 1.50) so now you have 1.5B = A . You are solving for B. (1.5 is the same as 3/2) solving for B you get B = 2A/3 . You are looking for how much less than A is B. so 1/3 more of B would yield A or B is 1/3 of A less. converting 1/3 to a percentage yields 33.33333 %